OpenStudy (anonymous):

How would I simplify the fraction: sqrt (w^4)/(x^3)?

OpenStudy (anonymous):

well the square root of w^4 would be w^2 (as w^2 * w^2 = w^4) so you would have w*sqrt(w^2)/(x^3)

OpenStudy (campbell_st):

well it can be written as \[\frac{\sqrt{w^4}}{\sqrt{x^3}} = \frac{\sqrt{w^2} \times \sqrt{w^2}}{\sqrt{x^2}\times \sqrt{x}}\] just simplify it... and use the rule \[\sqrt{a^2} = a\]

OpenStudy (anonymous):

keep in mind however sqrt(x^2) is |x|, not just x so after simplifying you would have |w|^2 although it really doesn't matter in this case because w is getting squared again anyways